An aqueous solution of HCl is added to an aqueous solution of Ca(OH) 2 and a reaction occurs. 8.9x10 -

6 mol of (OH – ) ions are measured in the 750mL solution. Remember, Molarity = mol/L. Calculate the

concentration of the H + ions. Is the solution more acidic or basic?


1
Expert's answer
2022-06-14T09:20:04-0400

CM (OH-) = n / V = 8,9*10-6 / 0,75 = 1,14*10-5


pOH = -log (OH-)

pOH = -log (1,14*10-5) = 4,94


pH + pOH = 14

pH = 14 - pOH = 14 - 4,94 = 9,06


pH = -log (H+)

9,06 = -log (H+)

H+ = 1,148*10-10 mol/l

Solution is basic

Answer:

[H+]=1,148*10-10

solution is a basic